Jim invested $24,000, part at 8%, and the rest at 7.2%. How much did he invest at each rate if his income from the 8% investment is two thirds that of the 7.2% investment?
Answer by josmiceli(6784)

let x be the amount invested at 8%
the remainder, $24,000 - x, is invested at 7.2%
the statement of the problem is:
.08x = 2/3 * (.072)(24,000 - x)
.08x = 2 * (.024)(24,000 - x)
.08x = (.048)(24,000 - x)
.08x + .048x = 1152
.128x = 1152
x = 9,000
24,000 - x = 15,000
check answers by plugging into first equation